OpenStudy (anonymous):

The length of a rectangle is 3 inches more than its width, and its perimeter is 22 inches. Find the width of the rectangle. a)4 in. b)5 in. c)7 in.

OpenStudy (oaktree):

So l+3=w, and 2l+2w=22, if l= length and w=width. Following?

OpenStudy (anonymous):

yeah

OpenStudy (oaktree):

So simplify 2l+2w=22 by dividing each term by 2 and get l+w=11

OpenStudy (oaktree):

Then subtract w from both sides and get l=11-w

OpenStudy (oaktree):

Now you know what l is, so just plug it into the other equation.

OpenStudy (oaktree):

You get (11-w)+3=w and simplify to get -w+14=w

OpenStudy (oaktree):

Simplify and get w=7

OpenStudy (oaktree):

So your answer is c, 7.

OpenStudy (anonymous):

ok thanks! :)